1. Show how to split the x term so that factorization can occur (Complete the square)
a. 2x2 +7x+3
b. 3x2 +5x–2
c. 4x2+11x+6
) d. 4x2 –6x–4
e. 8x2 +14x–4
f. 10x2 –19x–15

Answer by JellyBeans(8)

These are only my guesses but I'm pretty sure they're right.
a. 2x2 +7x+3 = (2x+1)(x+3)
b. 3x2 +5x–2 = (3x-1)(x+2)
c. 4x2+11x+6 = (4x+3)(x+2)
d. 4x2 –6x–4 = (2x-4)(2x+1)
e. 8x2+14x–4 = (8x-2)(x+2)
f. 10x2 –19x–15 = (5x+3)(2x-5)

I answered them just by staring at them in seconds